Given the following data, calculate the current five year interest rate: [Hint: There is more data...

  1. Given the following data, calculate the current five year interest rate: [Hint: There is more data here than you actually need.]

i1t

2.0%

i2t

2.8%

i1t+1

3.5%

i1t+2

3.0%

i1t+3

5.0%

i3t

2.8%

i1t+4

4.5%

i4t

3.4%

i1t+5

5.5%

I need help solving this with step by step instructions please!!
Answer #1

Answer

Formula :

int = (i1t + i1t+1 + i1t+2 + --------+ i1t+n-1)/n

where int = n period interest rate, i1t+n-1 = one year interest rate n years from now.

Using above data we have :

i5t = (i1t + i1t+1 + i1t+2 + i1t+3 + i1t+4)/5 and we have to calculate i5t

=> i5t = (2 + 3.5 + 3 + 5 + 3.4)/5 = 3.38 = 3.4(approx)

Hence, 5 years interest rate is 3.38 or 3.4(approx)